# Discovering reduced order model equations of many-body quantum systems using genetic programming: a technical report

### v1

@@ Abstract
In this technical report we present first results of applying Genetic Programming (GP) to obtain equations for constructing reduced order models of quantum systems, with a particular interest in nuclear Density Functional Theory. We employ the reduced basis method to obtain reduced coordinates as the amplitudes of the reduced basis, and use GP to avoid the need of constructing the reduced equations through, for example, a Galerkin projection. The reduced order models constructed through GP show excellent accuracy and speed performance, including extrapolations in the controlling parameters, and show promise as an effective method for emulating computationally demanding calculations in nuclear physics.
@@ I Introduction
Computational models for many-body quantum systems have quickly grown in complexity as theories get refined, more experimental data becomes available, and computers become more powerful. In fields with practical applications, such as nuclear physics, the demand for more accurate descriptions of nature, together with a push for predictions with well-quantified uncertainties [1, 2], has driven various efforts in the field to create surrogate models that can learn low-dimensional representations of the underlying system equations [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Several of the developed surrogate models are based on model reduction approaches [30, 31, 32] that approximates the solution manifold of equations parametrized by control variables $\alpha$ (for example, model parameters) by identifying reduced coordinates for the system, and then constructing equations for these coordinates.
In particular, the reduced basis method (RBM) [33, 34] traditionally works by identifying the reduced coordinates as the amplitudes of a reduced basis informed by previous high fidelity (HF) evaluations of the full system, and creates the reduced equations by projecting the underlying operators using the Petrov-Galerkin scheme [23]. This approach has shown great performance with speed ups of the original calculations usually between 2 and 7 orders of magnitude, yet it has been challenging to successfully apply it when the underlying operators are non-affine [28] or non-linear [25], usually requiring the mitigation of such features by hyperreduction schemes [35] such as the Empirical Interpolation Method [36, 37]. Alternative avenues to construct the reduced equations without explicitly projecting the original full system’ model, but rather by relying on data directly obtained from HF calculations have recently started to be explored [38, 39, 40, 41] (see for example [42] for a related application in other fields). Some of these developments follow a supervised machine learning (SML) approach in where various algorithms are trained to reproduce the dynamics of the reduced coordinates as the control parameters $\alpha$ change. In this work, we explore the application of Genetic Programming (GP) to obtain the reduced equations from data. This is achieved by creating an ensemble of initial possible equations that can reproduce the response of the reduced coordinates to parametric changes, and then performing an evolution algorithm in which the current population of models goes through changes to improve their fitness, measured as their accuracy in capturing the underlying reduced dynamics. We test this framework in two example problems [39] that exhibit non-affine and non-linear characteristics that make them problematic to emulate using the traditional Galerkin projection approach.
@@ II.1 Quantum systems
The first application problem we consider is a modified version of the Gross-Pitaevskii equation [43, 44] (a nonlinear Schrodinger equation used to describe dilute Bose-Einstein condensates):
@@
in where the Hamiltonian for each wave function $\phi^{(i)}$ consists of a trapping term proportional to $x^{2}$, as well as with a density $\rho(x)=\sum_{i}^{N}|\phi^{(i)}(x)|^{2}$ compromised by all $N$ particles in the system. The control parameters in this case are $\alpha=\{\kappa,q,\sigma\}$. Figure 1 shows the wave functions and their associated density for an example value of the control parameters, while Table 1 shows the ranges of the three parameters explored.
The second application problem we consider is Density Functional Theory (DFT) for a Skyrme type interaction [3]. DFT is used in nuclear physics to describe the atomic nucleus through the mean-field perspective, where the nucleons (protons and neutrons) interact with average densities and currents that are constructed from the nucleon wavefunctions (see  [45] for more details). The Hamiltonian of the system can be written as:
@@
where the wavefunction of each orbital $\phi_{i}$ interacts with a Hamiltonian $\hat{h}_{\alpha}^{(i)}$ that depends on all of them $\Phi=\{\phi^{(i)}\}_{i=1}^{N}$. This single particle Hamiltonian is derived from the Skyrme effective interaction [46, 47, 48, 49], with the nuclear part is made of time-even densities [50]:
@@
with the subscript $t=(0,1)$ denoting isoscalar and isovector densities, respectively. The parameters of this EDF model the interaction between the wavefunctions and the respective nucleonic densities in question (the kinetic energy density. We usually re-parametrize these constants by terms representing nuclear-matter properties [51, 52]:
@@
For this work we focus on 48Ca which, within the spherical DFT construction contains 6 independent proton orbitals and 7 independent neutron orbitals. Table 1 shows the ranges of the two varied parameters for this study, $J_{\text{sym}}^{\text{NM}}$ and $L_{\text{sym}}^{\text{NM}}$, which represent the symmetry energy and its slope [52] respectively.
@@
Table 1: Parameters varied for each of the two explored problems with their associated ranges. For the DFT case, both the symmetry energy $J_{\text{sym}}^{\text{NM}}$ and its slope $L_{\text{sym}}^{\text{NM}}$ are expressed in MeV. A total of 492 evaluations of the modified Gross-Pitaevskii densities were used for training and 212 for testing. A total of 25 evaluations of the DFT protons and neutrons densities were used for training and 16 for testing.
@@ II.2 Reduced basis method
where we rewrite the equation in terms of the operator $F_{\alpha}$ which is, for example $F_{\alpha}[\phi^{(i)}]=\big{[}\hat{h}_{\alpha}^{(i)}[\Phi]-E_{i}\big{]}\phi^{(i)}$ in the DFT case 2. For each of the two examples described (1) and (2) we could approximate each wave functions as:
@@
where the few $n$ coefficients $a_{k}(\alpha)$ represent latent reduced coordinates that depend on the control variables $\alpha$, while the reduced basis $\phi_{k}^{(i)}$ is constructed informed by previous HF solutions for different values of $\alpha$. In several recent successful RBM applications to nuclear physics [24, 23, 25, 28, 41, 29], the reduced basis is constructed by performing a principal component analysis [53], or singular value decomposition [54] on the set of HF solutions.
Figure 2: Densities calculated for various values of the controlling parameters $\alpha$ specified in Table 1. The inset plots show the exponential decay of the singular values from the SVD. Panel a) Density $\rho$ for the modified Gross-Pitaevskii Eq. (1). Panel b) Densities of neutrons $\rho_{n}$ and decay of singular values for both protons $\rho_{p}$ and neutrons.
The second step is to obtain the equations that describe the response of the latent variables $a_{k}$ to changes in the parameters $\alpha$. In many RBM applications (including several of the previously cited work), a way to achieve this is to project the equations into a subspace spanned by test functions $\psi_{j}$ [23, 33, 34]:
@@
In our previous work [23], we successfully constructed a reduced order model for the Gross-Pitaevskii equation (1) in the case of $N=1$ and $\sigma=1$ using the described Galerkin projection scheme (7), yet for bigger powers $\sigma$ for the density, or even fractional powers, this approach proves inadequate to create efficient reduced equations. The challenge lies in the inability to pre-compute the Galerkin projections in the offline stage of building the emulator, since the equations are non-affine and non-linear in $\sigma$ and the wavefunctions. Various terms in realistic EDFs [55, 56, 57] contains fractional powers of densities, which motivates the development of alternative approaches to obtain reduced equations, such as the one we present next.
@@ II.3 Genetic Programming
Genetic Programming (GP)  [^1] Often it is abbreviated as GP, but not to be confused with the same abbreviation commonly used in nuclear physics emulation for Gaussian Processes. Instead, it is a bio-inspired search method making use of processes of stochastic variation and cumulative selection. is a type of Artificial Intelligence (AI) that falls under the broader category of evolutionary computation (EC), which is a subset of AI inspired by biological evolution mechanisms such as selection, mutation, and crossover to solve problems [58, 59, 60, 61]. Specifically, GP is a population-based stochastic iterative search algorithm in the search space of computer programs. Like other evolutionary meta heuristics, GP evolves a set of candidate solutions (the population) by mimicking the basic principles of Darwinian evolution. The evolutionary process involves an iterative application of a fitness-based selection of the candidate solutions and their variation throughout genetically-inspired operators, such as crossover and mutation [58]. Three of the most typical representations employed in GP are expression trees in tree GP (TGP), linear sequences of instructions in linear GP (LGP) and circuit-type graphs in Cartesian GP (CGP). Here we shall be mostly concerned with TGP, and want to only briefly touch the other two representations. In this work, we will use TGP and will refer to it as simply GP. In GP, the evolving programs are constructed by composing elements belonging to two specific, predefined, sets: a set of primitive functions $F$, which appear as the internal nodes of the trees, a set of terminal input features $T$ and constant values $C$, which represent the leaves of the trees. In the context of SML problem-solving, the trees encode symbolic expressions mapping inputs $X$ to the outputs $y$. Typically, GP is used with the so-called subtree mutation and swap crossover [58]. The latter exchanges two randomly selected subtrees between two different parent individuals. The former randomly selects a subtree in the structure of the parent individual and replaces it with a new, randomly generated tree. Algorithm 1 provides a pseudo-code for the GP, whereas Figure 3 provides a visual representation of GP-models and their mapping between tree and function.
@@
In an ML context, GP provides a user with a completely different experience in terms of transparency and interpretability of the final models when compared to more “black-box” models of, say, deep artificial neural networks (ANNs). This is because GP typically outputs symbolic expressions or tree structures representing programs, which are generally interpretable and can be analyzed by humans. By manipulating discrete structures, the inner workings of a GP-based model are clearer, and the relationships between features and predictions are explicitly defined. As a result, evolved solutions can offer valuable insights into how the model arrives at its predictions, allowing humans to understand the decision-making process and relate this to their domain-knowledge.
@@
In both cases we select a total of three basis $n=3$. For the DFT case we analyze the proton density $\rho_{p}$ as well as the neutron density $\rho_{n}$. The reduced basis $\rho_{k}(x)$ for each of the respective densities is obtained following a singular value decomposition on a set of HF solutions for the training parameters (see Table 1). Note that, since the original equations Eq. (1) and Eq. (2) can’t directly be solved by only considering the density (they involve all wave functions), this GP approach is effectively learning reduced equations that would not be possible to obtain through a direct Galerkin projection.
Table 2 lists the hyper-parameters (HPs) used for GP, along with cross-validation settings. The HPs were selected following common practices found across the literature to avoid a computationally demanding tuning phase. $R^{2}$ was used as a fitness function [90, 91, 92] as it was found to converge faster, generalize better, even when only a few data points are available. A common practice in GP is to protect operators to avoid undefined mathematical behaviour of the function set by defining some ad-hoc behaviour at those points like, for instance, returning the value 1 in the case of a division by zero to make it possible for genetic programming to synthesize constants by using $x/x$ [58]. However, it was shown that these techniques present several shortcomings in the vicinity of mathematical singularities [90]. In this study, programs that produced invalid values were automatically assigned a bad fitness value, making them unlikely to be selected. By diminishing the selection attractiveness of solutions that produced invalid values, we expect to obtain models whose fitness landscape is less sharp [93].
@@
To contrast the GP ability to learn the equations from data, we also implemented a linear regression model. In the context of SML, linear models are a class of algorithms that assume a linear relationship between the input variables (features) and the output variable (target), in this case the controlling parameters $\alpha$ and the reduced coefficients $a_{k}$, respectively. To make a prediction, the input features are multiplied by the learnable weights, summed together and added a learnable constant called the intercept or bias [96]. Formally: $y=\beta_{0}+\beta_{1}x_{1}+\beta_{2}x_{2}+\ldots+\beta_{n}x_{n}$, where $y$ is the target, $\beta_{0}$ is the intercept, and $\beta_{k}$ is a learnable weight associated with input feature $x_{k}$. The learnable weights and the intercept are usually estimated from the data using Ordinary Least Squares (OLS) estimation procedure to minimize the sum of squared residuals (i.e., difference between the observed target values in the training dataset, and the values predicted by the linear model). Formally: $\min_{\beta}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$, where $\hat{y}_{i}$ is the model’s prediction for data instance $i$.
@@
Figure 4: Performance analysis of the SML approaches for the modified Gross-Pitaevskii Eq. (1). Panel a): Computational Accuracy vs Time plot [28] that shows for each of the 212 testing evaluations the norm of the residual in the vertical axis and the calculation time in the horizontal axis for cubic linear regression with L1 regularization (dark red), with L2 regularization (dark green), and with GP (blue). The corresponding dashed vertical lines are the median time for each method, while the dotted horizontal lines the median error. The horizontal dashed black line shows the median error obtained when optimizing the coefficients in Eq. (8) directly against the test solution. The vertical grey solid line shows the median calculation time of the HF solver Python implementation, which is more than 4 orders of magnitude slower than any of the three methods. Panels b) and c) show the evolution of the first coefficient of the reduced basis $a_{1}$ as the parameters $\kappa$ and $\sigma$ change, respectively, while the other parameters are held constant. The black dashed line shows the optimal value of the coefficient as the parameters change. The dashed vertical grey line shows the boundary on which the SML methods were trained, after which the calculations are effectively extrapolating.
@@
In order to constraint the complexity of the model and improve generalization ability, regularization terms can be added to the sum of squared residuals. L1 regularization promotes models’ sparsity and is useful to tackle high dimensional tasks (i.e. with many input features), especially if many are expected to be irrelevant. L2 regularization is frequently used when the input features are expected to be informative but their impact has to be reduced to avoid overfitting, and/or are highly correlated (multicollinear). Formally, the L1 term is given by $\lambda_{1}\sum_{k=0}^{n}|\beta_{k}|$, whereas L2 term is given by $\lambda_{2}\sum_{k=0}^{n}\beta_{k}^{2}$. For creating high-order interactions between the original input features $X$, one can generate a new feature set $X^{\prime}$ consisting of all polynomial combinations of the features up to a given degree. Provided $X^{\prime}$, a linear model can learn non-linear relationships. In this study, we use regularized linear regression with polynomial terms and we rely on scikit-learn implementation [97, 98].
@@
Note that regularized linear regression has been long used in scientific applications. Sparse-identification of non-linear dynamics (SINDy) is a popular method for discovering governing equations in dynamic systems [31]. Essentially it utilizes linear regression with the L1 regularizing term to promote sparsity, yielding smaller models, which can offer better interpretability, which is considered an appealing property. SINDy has been successfully applied to find governing equations in many different applications [99, 100, 101, 102, 103]. Nevertheless, SINDy-like approaches present a conspicuous disadvantage: all of the non-linear relationships to be included in the model search must be manually pre-defined, which requires some degree of expertise. This can be particularly limiting when applied to a problem where valuable non-linear relationships are unexpected, and is one of the advantages we expect from GP.

### v2

# Discovering Reduced-order Model Equations of Many-body Quantum Systems using Genetic Programming: A Technical Report

@@ Abstract
Here we report first results of applying Genetic Programming to obtain equations for constructing reduced-order models of quantum systems, with a particular interest in nuclear Density Functional Theory. We employ the reduced basis method to obtain reduced coordinates as the amplitudes of the reduced basis, and use genetic programming to avoid the need of constructing the reduced equations through, for example, a Galerkin projection. The resulting reduced-order models show excellent performance in both accuracy and speed, including for extrapolations in the control parameters, and show promise as an effective method for emulating computationally demanding calculations in Nuclear Physics.
@@ I Introduction
Computational models for many-body quantum systems have quickly grown in complexity as theories were refined, more experimental data became available, and computers grew more powerful. In fields with practical applications, such as nuclear physics, the demand for more accurate descriptions of nature, together with a push for predictions with well-quantified uncertainties [1, 2], has driven various efforts to create surrogate models that can learn low-dimensional representations of the underlying system equations [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Several of the developed surrogate models are based on model reduction approaches [30, 31, 32] that approximate the solution manifold of equations parametrized by control variables $\alpha$ (for example, model parameters) through identification of reduced coordinates for the system, followed by the construction of equations for these coordinates.
In particular, the reduced basis method (RBM) [33, 34] traditionally works by identifying reduced coordinates as the amplitudes of a reduced basis informed by previous high fidelity (HF) evaluations of the full system. The reduced equations are created by projecting the underlying operators using the Petrov-Galerkin scheme [23]. This approach has shown good performance with speed-ups compared to the original calculations of between 2 and 7 orders of magnitude. However, it has been challenging to successfully apply this method when the underlying operators are non-affine [28] or non-linear [25], usually requiring the mitigation of such features by hyperreduction schemes [35] such as the Empirical Interpolation Method [36, 37]. Alternative avenues to construct reduced equations without explicitly projecting the original full system model, but rather by relying on data directly obtained from HF calculations have recently started to be explored [38, 39, 40, 41] (see for example [42] for a related application in other fields). Some of these developments follow a supervised machine learning (SML) approach whereby various algorithms are trained to reproduce the dynamics of the reduced coordinates as control parameters $\alpha$ change. In this work, we explore the application of Genetic Programming (GP) to obtain the reduced equations from data. This is achieved by creating an ensemble of initial possible equations that can reproduce the response of the reduced coordinates to parametric changes, and then executing an evolutionary algorithm in which the current population of models undergoes changes to improve their fitness, measured as their accuracy in capturing the underlying reduced dynamics. We test this framework in two example problems [39] which exhibit non-affine and non-linear characteristics that make them problematic for emulation under the traditional Galerkin projection approach.
@@ II.1 Quantum Systems
The first application problem we consider is a modified version of the Gross-Pitaevskii equation [43, 44] (a nonlinear Schrödinger equation used to describe dilute Bose-Einstein condensates):
@@
Here, the Hamiltonian for each wave function $\phi^{(i)}$ consists of a trapping term proportional to $x^{2}$, as well as a term depending on the density $\rho(x)=\sum_{i}^{N}|\phi^{(i)}(x)|^{2}$ of all $N$ particles of the system. The control parameters in this case are $\alpha=\{\kappa,q,\sigma\}$. Figure 1 shows the wave functions and their associated density for an example value of the control parameters, while Table 1 shows the ranges of the three parameters explored.
The second application problem we consider is based on Density Functional Theory (DFT) for a Skyrme type interaction [3]. DFT is used in nuclear physics to describe the atomic nucleus in a mean-field perspective, where the nucleons (protons and neutrons) interact with average densities and currents that are constructed from the nucleon wavefunctions (see  [45] for more details).
The Hamiltonian of the system can be written as:
@@
where the wavefunction of each orbital $\phi_{i}$ interacts with a Hamiltonian $\hat{h}_{\alpha}^{(i)}$ that depends on all of them $\Phi=\{\phi^{(i)}\}_{i=1}^{N}$. This single particle Hamiltonian is derived from the Skyrme effective interaction [46, 47, 48, 49], with the nuclear part made of time-even densities [50]:
@@
with the subscript $t=(0,1)$ denoting isoscalar and isovector densities, respectively. The parameters of this Energy Density Functional (EDF) model the interactions between the wavefunctions and the respective nucleonic densities in question (the kinetic energy density). We usually re-parametrize these constants by terms representing nuclear-matter properties [51, 52]:
@@
For this work we focus on 48Ca which, contains 6 independent proton orbitals and 7 independent neutron orbitals within the spherical DFT construction. Table 1 shows the ranges of the two parameters varied for this study, $J_{\text{sym}}^{\text{NM}}$ and $L_{\text{sym}}^{\text{NM}}$, which represent the symmetry energy and its slope [52] respectively.
@@
Table 1: Parameters varied for each of the two problems explored, with their associated ranges. For the DFT case, both the symmetry energy $J_{\text{sym}}^{\text{NM}}$ and its slope $L_{\text{sym}}^{\text{NM}}$ are expressed in MeV. A total of 492 evaluations of the modified Gross-Pitaevskii densities were used for training and 212 for testing. A total of 25 evaluations of the DFT proton and neutron densities were used for training and 16 for testing.
@@ II.2 Reduced Basis Method
where we rewrite the equation in terms of the operator $F_{\alpha}$ which is, for example $F_{\alpha}[\phi^{(i)}]=\big{[}\hat{h}_{\alpha}^{(i)}[\Phi]-E_{i}\big{]}\phi^{(i)}$ in the DFT case 2. For each of the twoboth examples described (1) and (2) we could approximate each wave functions as:
@@
where an assumed small number $n$ of coefficients $a_{k}(\alpha)$ represent latent reduced coordinates that depend on the control variables $\alpha$, while the reduced basis $\phi_{k}^{(i)}$ is constructed informed by previous HF solutions for different values of $\alpha$. In several recent successful RBM applications to nuclear physics [24, 23, 25, 28, 41, 29], the reduced basis is constructed by performing a principal component analysis [53], or singular value decomposition [54] on the set of HF solutions.
Figure 2: Densities calculated for various values of the controlling parameters $\alpha$ specified in Table 1. Inset plots show the exponential decay of singular values from the SVD. (a) Density $\rho$ for the modified Gross-Pitaevskii Eq. (1). (b) Densities of neutrons $\rho_{n}$ and decay of singular values for both protons $\rho_{p}$ and neutrons.
The second step is to obtain equations that describe the response of the latent variables $a_{k}$ to changes in the parameters $\alpha$. In many RBM applications (including several of the previously cited works), a way to achieve this is to project the equations into a subspace spanned by test functions $\psi_{j}$ [23, 33, 34]:
@@
In our previous work, [23], we successfully constructed a reduced-order model for the Gross-Pitaevskii equation (1) in the case of $N=1$ and $\sigma=1$ using the described Galerkin projection scheme (7), yet for higher powers $\sigma$ of the density, or even for fractional powers, this approach proves inadequate to create efficient reduced equations. The challenge lies in the inability to pre-compute the Galerkin projections in the offline stage of building a emulator, since the equations are non-affine and non-linear in $\sigma$ and the wavefunctions. Various terms in realistic EDFs [55, 56, 57] contain fractional powers of densities, which motivates the development of alternative approaches to obtain reduced equations, such as the one we present next.
@@ II.3 Genetic Programming
Genetic Programming (GP)  [^1] Often it is abbreviated as GP, but not to be confused with the same abbreviation commonly used in nuclear physics emulation for Gaussian Processes. Instead, it is a bio-inspired search method making use of processes of stochastic variation and cumulative selection. is a type of Artificial Intelligence (AI) that falls under the category of evolutionary computation (EC), which is a subset of AI inspired by biological evolution mechanisms such as selection, mutation, and crossover to solve problems [58, 59, 60, 61]. Specifically, GP is a population-based stochastic iterative search algorithm in the search space of algorithms or computer programs. Like other evolutionary meta-heuristics, GP evolves a set of candidate solutions (the population) by mimicking the basic principles of Darwinian evolution. The evolutionary process involves an iterative application of a fitness-based selection of candidate solutions and their variation by genetically-inspired operators such as crossover and mutation [58]. Three of the most typical representations employed in GP are expression trees in tree GP (TGP), linear sequences of instructions in linear GP (LGP) and circuit-type graphs in Cartesian GP (CGP). Here we shall be mostly concerned with TGP, and want to only briefly touch the other two representations. In this work, we will use TGP and will refer to it as simply GP. In GP, the evolving programs are constructed by composing elements belonging to two specific, predefined, sets: a set of primitive functions $F$, which appear as the internal nodes of the trees, and a set of terminal input features $T$ and constant values $C$, which represent the leaves of the trees. In the context of supervised machine learning problems, the trees encode symbolic expressions mapping inputs $X$ to outputs $y$. Typically, GP is used with the so-called subtree mutation and swap crossover [58]. The latter exchanges two randomly selected subtrees between two different parent individuals. The former randomly selects a subtree in the structure of the parent individual and replaces it with a new, randomly generated tree. Algorithm 1 provides a pseudo-code for the GP, whereas Figure 3 shows a visual representation of GP-models and their mapping between tree and function.
@@
In an ML context, GP provides a user with a completely different experience in terms of transparency and interpretability of the final models when compared to more “black-box” models of, say, deep artificial neural networks (ANNs). This is because GP typically outputs symbolic expressions or structures representing programs, which are generally interpretable and can be analyzed by humans. By manipulating discrete structures, the inner workings of a GP-based model are clearer, and the relationships between features and predictions are explicitly defined. As a result, evolved solutions can offer valuable insights into how the model arrives at its predictions, allowing humans to understand the algorithmic decisions and relate this to their domain-knowledge.
@@ II.4 Generating Reduced-order Equations with GP
In both cases we select a total of three basis components $n=3$. For the DFT case we analyze the proton density $\rho_{p}$ as well as the neutron density $\rho_{n}$. The reduced basis $\rho_{k}(x)$ for each of the respective densities is obtained following a singular value decomposition on a set of HF solutions for the training parameters (see Table 1). Note that, since the original equations Eq. (1) and Eq. (2) can’t directly be solved by only considering the density (they involve all wave functions), this GP approach is effectively learning reduced equations that would not be possible to obtain through a direct Galerkin projection.
Table 2 lists the hyper-parameters (HPs) used for GP, along with cross-validation settings. The HPs were selected following common practice across the literature in order to avoid a computationally demanding tuning phase. $R^{2}$ was used as fitness function [90, 91, 92] since it was found to converge faster and generalize better, even when only a few data points are available. A common practice in GP is to protect operators from undefined mathematical behaviour through defining some ad-hoc behaviour at those points like, for instance, returning the value 1 in the case of a potential division by zero to make it possible for genetic programming to synthesize constants by using $x/x$ [58]. However, it was shown that these techniques present several shortcomings in the vicinity of mathematical singularities [90]. In this study, programs that produced invalid values were automatically assigned a low fitness values, making them unlikely to be selected. By reducing the selection probability of solutions that produce invalid values, we expect to obtain models whose fitness landscape is less ”sharp” [93].
@@
To highlight the ability of genetic programming to learn the equations from data, we also implemented a linear regression model. In the context of SML, linear models are a class of algorithms that assume a linear relationship between the input variables (features) and the output variable (target), in this case the controlling parameters $\alpha$ and the reduced coefficients $a_{k}$, respectively. To make a prediction, the input features are multiplied by the learnable weights, summed together and added a learnable constant called the intercept or bias [96]. Formally: $y=\beta_{0}+\beta_{1}x_{1}+\beta_{2}x_{2}+\ldots+\beta_{n}x_{n}$, where $y$ is the target, $\beta_{0}$ is the intercept, and $\beta_{k}$ is a learnable weight associated with input feature $x_{k}$. The learnable weights and the intercept are usually estimated from the data using Ordinary Least Squares (OLS) estimation procedure to minimize the sum of squared residuals (i.e., difference between the observed target values in the training dataset, and the values predicted by the linear model). Formally: $\min_{\beta}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$, where $\hat{y}_{i}$ is the model’s prediction for data instance $i$.
@@
Figure 4: Performance analysis of supervised machine learning approaches for the modified Gross-Pitaevskii Eq. (1). (a): Computational Accuracy vs Time plot [28] that shows for each of the 212 test evaluations the norm of the residual on the vertical axis and the calculation time on the horizontal axis for cubic linear regression with L1 regularization (dark red), with L2 regularization (dark green), and with GP (blue). Dashed vertical lines are the median time for each method, while dotted horizontal lines depict the median error. The horizontal dashed black line shows the median error obtained when optimizing the coefficients in Eq. (8) directly against the test solution. The vertical grey solid line depicts the median calculation time of the HF solver in Python implementation, which is more than 4 orders of magnitude slower than any of the three methods. (b) and (c) show the evolution of the first coefficient of the reduced basis $a_{1}$ as parameters $\kappa$ and $\sigma$ change, while the other parameters are held constant. The black dashed line shows the optimal value of the coefficient as the parameters change. The dashed vertical grey line represents the training boundary, beyond which the calculations are effectively extrapolating.
@@
In order to constrain the complexity of models and improve their generalization ability, regularization terms can be added to the sum of squared residuals. L1 regularization promotes model sparsity and is useful in high dimensional tasks (i.e. when there are many input features), especially if many are expected to be irrelevant. L2 regularization is frequently used when the input features are expected to be informative but their impact has to be reduced to avoid overfitting, and/or if they are highly correlated (multicollinear). Formally, a L1 term is given by $\lambda_{1}\sum_{k=0}^{n}|\beta_{k}|$, whereas a L2 term is given by $\lambda_{2}\sum_{k=0}^{n}\beta_{k}^{2}$. For considering higher-order interactions between the original input features $X$, one can generate a new feature set $X^{\prime}$ consisting of all polynomial combinations of the features up to a given degree. Provided such $X^{\prime}$, a linear model can learn non-linear relationships. In this study, we use regularized linear regression with polynomial terms and we use the scikit-learn implementation [97, 98].
@@
Note that regularized linear regression has long been used in scientific applications. Sparse-identification of non-linear dynamics (SINDy) is a popular method for discovering governing equations in dynamic systems [31]. Essentially it utilizes linear regression with the L1 regularizing term to promote sparsity, yielding smaller models, which can offer better interpretability, considered an appealing property. SINDy has been successfully applied to find governing equations in many different applications [99, 100, 101, 102, 103]. Nevertheless, SINDy-like approaches present a obvious disadvantage: Non-linear relationships are combinatorial in input variables or all of the non-linear relationships to be included in the model search must be manually pre-defined, requiring some degree of expertise. This can be particularly limiting when applied to problems where valuable non-linear relationships are unexpected, and is one of the advantages we expect from GP.